can you please help me find x?
2X10^-5 = (1+0.23x)^-3
Thank you in advance.
: can you please help me find x?
2X10^-5 = (1+0.23x)^-3
Thank you in advance.

Answer by josmiceli(2162)

2*10^(-5) = (1 + .23x)^(-3)
multiply both sides by this expression:
(10^5)*(1 + .23x)^3
(2*10^(-5))*(10^5)*(1 + .23x)^3) = ((1 + .23x)^(-3))*(1 + .23x)^3)*(10^5)
This reduces to
2*(1 + .23x)^3 = 10^5
(1 + .23x)^3 = .5*10^5
Take the cube root of both sides
1 + .23x = (50000)^(1/3)
1 + .23x = 36.84
.23x = 36.84 - 1
.23x = 35.84
x = 155.83 answer
Hope I got it right